# LMME3DHF: Benchmarking and Evaluating Multimodal 3D Human Face Generation with LMMs

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 27 - 31, 2025; Dublin, Ireland.isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Multimedia databasesccs: Information systems Multimedia streamingccs: Information systems Multimedia content creation
@@ 1. Introduction
In this paper, we introduce Gen3DHF, a comprehensive dataset and benchmark comprising 2,000 diverse 3D HF video samples generated by five distinct models. As illustrated in Figure 1, we collect 90,000 human annotations, independently assessing quality and authenticity, resulting in 4,000 Mean Opinion Scores (MOS). Additionally, we also collect the distortion markings and corresponding distortion descriptions, resulting in a total of 2,000 distortion-aware saliency maps and corresponding distortion descriptions. Based on Gen3DHF, we propose LMME3DHF, a LMM-based metric specifically designed to not only evaluate 3D HF content across the two dimensions, i.e., quality and authenticity, but also to predict and output the salient distortion regions along with their corresponding textual descriptions. LMME3DHF leverages instruction tuning (Liu et al., 2023b) and LoRA adaptation(Hu et al., 2022) techniques to fine-tune the language model. Extensive experimental results demonstrate that LMME3DHF achieves state-of-the-art performance, surpassing existing quality assessment methods not only in accurately predicting quality scores for AI-generated 3D HF but also in effectively identifying distortion-aware salient regions and giving distortion types.
@@
  • Extensive experiments validate LMME3DHF’s superior performance in both accurately predicting quality scores for AI-generated 3D human faces and effectively identifying distortion-aware salient regions and distortion types, establishing it as a new state-of-the-art benchmark for evaluating AI-generated 3D human faces.
@@ 2. RELATED WORK
Recent advancements in 3D-aware generative adversarial networks (GANs) significantly improve the quality of synthesized human faces, achieving breakthroughs in photorealism, multi-view consistency, and fine-grained controllability. While explicit and implicit neural rendering methods each have limitations in scalability or fidelity, (Chan et al., 2022) introduced a hybrid explicit–implicit tri-plane architecture that achieves high-resolution, multi-view-consistent synthesis from unstructured 2D images. Building on it, (An et al., 2023; Sun et al., 2023b; Wu et al., 2022b, 2023b; Deng et al., 2024b, a; Or-El et al., 2022; Schwarz et al., 2022; Zhang et al., 2022; Sun et al., 2022b) extend 3D HF generation to full $360^{\circ}$ head views, introduce animatable controls over facial expressions and head–shoulder movements respectively, improvements in fine-grained 3D control, detailed facial deformation with high fidelity, demonstrated high-fidelity 4D head synthesis and motion transfer using synthetic or pseudo multi-view data, achieving robust motion control without relying on monocular 3D Morphable Model (3DMM) reconstruction. Nevertheless, 3D human face generation still suffers from certain distortions, such as inconsistent views across frames, unrealistic geometry in unseen regions, or artifacts introduced by limited training data and view ambiguity.
@@ 2.2. Video Quality Assessment
With the rapid advancement of AI-generated image (AIGI), various image quality assessment (IQA) methods (Li et al., 2023b; Wang et al., 2023; Liang et al., 2024; Kirstain et al., 2023; Wu et al., 2023a) have emerged to evaluate the perceptual quality of these images. As the capabilities of AI-generated content continue to evolve (Dhariwal and Nichol, 2021; Rombach et al., 2022), AI-generated videos (AGVs) are now receiving increasing attention and being applied across a wide range of use cases. To assess their quality, several video quality assessment (VQA) studies have been conducted (Chen et al., 2024a; Chivileva et al., 2023; Liu et al., 2024, 2023a; Kou et al., 2024; Huang et al., 2024a). Early video quality assessment (VQA) methods often adopted evaluation metrics originally designed for image quality assessment (IQA), such as the Inception Score (IS) (Salimans et al., 2016) and Fréchet Inception Distance (FID) (Heusel et al., 2017). While effective for measuring basic perceptual similarity, these metrics do not fully capture temporal coherence or motion consistency in videos. More recently, Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021) has been employed as a backbone for quality assessment tasks, leveraging its powerful multimodal representations to better align generated content with human-perceived semantics. Meanwhile, the rise of video large multimodal models introduces semantic-aware quality assessment, where models pre-trained on vision-language tasks jointly analyze visual fidelity and textual descriptions and particularly excel in visual question-answering. Existing deep learning-based VQA methods lack interactive QA capabilities, while video large multimodal models show limited precision in perceptual quality assessment.
@@ 2.3. Saliency Prediction
Recent advances in computer vision have driven significant progress in human visual attention analysis, with saliency prediction emerging as a particularly crucial research area. This technology simulates the human visual system’s mechanisms for allocating attention across different image regions. In recent years, numerous research efforts have focused on developing saliency prediction models. Traditional models (Bruce and Tsotsos, 2005; Goferman et al., 2011; Erdem and Erdem, 2013) primarily rely on low-level features such as intensity, color, and contrast to generate saliency maps. In contrast, deep learning-based models (Huang et al., 2015; Che et al., 2019; Lou et al., 2022) are capable of capturing more complex visual patterns and semantic information, allowing them to outperform traditional approaches across a range of tasks. However, despite their success in general applications, these models often fall short in the context of quality assessment, where identifying distortion-specific regions is crucial for accurately evaluating perceptual quality. This motivates us to develop a saliency prediction model specifically tailored for identifying distortion-aware salient regions in AI-generated human faces as distortion-aware saliency prediction not only guides model attention toward visually critical regions but also enhances interpretability by highlighting the exact areas that degrade user experience.
@@ 3.2. Subjective Experiment
To better align VQA with human preferences, AI-generated content should be evaluated from multiple perspectives. Therefore, this paper proposes a dual-perspective evaluation framework, focusing on quality and authenticity. The quality dimension assesses perceptual attributes such as texture, color fidelity, and structural integrity. In contrast, the authenticity dimension evaluates perceptual aspects like unidentified object appearance, head structure and the presence of unrealistic textures or shapes.
Figure 4. The subjective experiment user interface
Subjective experiments are conducted in accordance with the ITU-R BT.500-13 (Series, 2012) guidelines. The testing environment is an indoor setting with standardized lighting conditions, and participants are seated approximately 30 cm apart to minimize mutual interference. The experiment interface is display on a monitor with a resolution of 1920 $\times$ 1080 pixels. As shown in Figure 4, the interface presents a video along with a static image containing three snapshots from the video at $-45^{\circ}$, $0^{\circ}$ and $45^{\circ}$ angles (with $0^{\circ}$ representing the frontal view). The interface also includes interactive buttons “Previous”, “Replay” and “Next” to allow participants to navigate through the projection videos of 3D HF content. Additionally, two sliders ranging from 0 to 5 enable participants to score the content on both quality and authenticity dimensions smoothly. A “MARK” button is provided to allow participants to highlight any visible distortions on the static images. If distortions are marked, participants are instructed to provide a short, concise description of the issue. This can be done by typing manually using the keyboard or by recording their explanation through a “Voice” button. Audio that has been recorded will then automatically convert to words. A total of 24 subjects participate in the subjective experiment. All participants receive a thorough briefing beforehand. Each session lasts approximately two hours. Finally, we obtain a total of 90,000 human annotations including 84,000 reliable score ratings (21 annotators $\times$ 2 dimensions $\times$ 2,000 videos), 6,000 distortion-aware saliency maps and distortion descriptions (3 annotators $\times$ 2,000 images) which were then aggregated into a single combined annotation per image.
@@ 3.3. Subjective Data Processing
   (a)                                     (b)
Additionally, to quantify performance and standardize raw distortion descriptions, we classify distortions into nine distinct categories based on their visual characteristics $d_{i}$, where $\{d_{i}|_{i=1}^{9}\}=$”Eye Distortions”, ”Mouth Distortions”, ”Hair Distortions”, ”Facial Feature Distortions”, ”Head Structure Distortions”, ”Overlap or Blending Issues”, ”Blurring / Exposure / Grain”, ”Accessories or Cloth Distortions”, “No Distortion”}. Due to the nature of our subjective experiment data where human annotators marke distortion regions using red dots, a transformation is required to convert these discrete annotations into continuous distortion-aware saliency maps. As illustrated in Figure 5, and following standard practices in saliency prediction tasks, we first identify the center of each distorted region based on the human-marked red dots. These fixation maps are then smoothed using a Gaussian filter with a standard deviation of $\sigma=5$.
@@ 3.4. Subjective Data Analysis
The MOS and score distributions for the quality and authenticity dimensions are illustrated in Figure 2(a). Both distributions are slightly right-skewed, indicating that most samples received moderate to high ratings, peaking around the 55–60 range. This suggests that the majority of the generated 3D human face (HF) content was perceived as reasonably convincing in terms of both visual quality and authenticity, while only a small portion received lower ratings. To further investigate the perceptual performance of different 3D HF generation models, we compare their MOS for quality and authenticity. As shown in Figure 2(b), EG3D (Chan et al., 2022) performs well in terms of quality but lags in authenticity, whereas PanoHead (An et al., 2023) exhibits the opposite trend, achieving higher authenticity scores but slightly lower quality ratings. This contrast underscores the importance of evaluating quality and authenticity as separate dimensions, as a single unified metric may fail to capture the nuanced perceptual differences between models. As illustrated in Figure 3(a), each distortion type occurs regularly, with facial feature distortions appearing most frequently. This indicates that AI-generated 3D HFs particular challenges in generating smooth, realistic, and detailed facial features. The distribution of distortion types across different generation models is illustrated in Figure 3(b). As shown in Figure 6, we observe that an increased number of distortion-aware salient regions strongly correlates with a significantly worse human visual experience. In other words, lower scores are associated with more extensive and concentrated saliency in distorted areas.
Figure 7. Overview of the LMME3DHF architecture. The model includes three functions: (a) text-defined quality level prediction, (b) precises score prediction, (c) distortion-aware saliency map prediction. The training process consists of three stages: (a) instruction tuning of the model via text-defined levels and distortion description, (b) fine-tuning the LLM via numerical scores, (c) train saliency decoder from scratch.
@@ 4.1. Model Structure
Visual Encoding.  As shown in Figure 7, we leverage the pretrained Vision Transformer (ViT) module embedded in Qwen2.5-VL-7B to extract visual features. This model demonstrates strong capabilities in handling videos with varying resolutions and frame rates, allowing us to bypass typical preprocessing steps such as compression or resizing, which often lead to information loss critical for video quality evaluation. The native projector, implemented as a two-layer Multi-Layer Perceptron (MLP), is used to project the extracted visual features into a latent space aligned with the text embeddings used in the language model.
@@ 4.2. Training and Fine-tuning Strategy
Instruction Tuning.  Due to the inherent strengths of LLMs in processing textual rather than numerical data, we begin by converting continuous quality scores into categorical, text-based quality levels. To mitigate the issue of data imbalance caused by the clustering of MOS values within a narrow range, we uniformly divides the score range from the minimum value (m) to the maximum value (M) into five equal intervals and assigns each interval a corresponding textual label based on Equation (4).
@@
where $\{l_{i}|_{i=1}^{5}\}=\{bad,poor,fair,good,excellent\}$ that standard text rating levels defined by ITU (Series, 2012). The distortion descriptions were categorized into nine classes, as previousl literature described. During instruction tuning, the model is trained using language loss, guiding it to generate quality-level predictions and distortion category classifications.
Quality Regression Fine-Tuning.  To enable continuous quality score prediction, we implement a quality regression module. This module takes the last hidden states from the LLM as input and outputs precise predicted scores. As full fine-tuning of LLMs can be computationally expensive, we leverage the LoRA (Hu et al., 2022) technique to efficiently adapt the model to the quality regression task. During this fine-tuning phase, we use the L1 loss function to minimize the difference between the predicted scores and the ground-truth MOS, ensuring accurate and reliable quality assessment.
@@ 5.1. Experiment Setup
To evaluate the correlation between predicted scores and the ground-truth MOS, we employ three evaluation metrics: Spearman’s Rank Correlation Coefficient (SRCC), Pearson’s Linear Correlation Coefficient (PLCC), and Kendall’s Rank Correlation Coefficient (KRCC). For assessing visual question answering performance, we use average accuracy as the evaluation metric. For the distortion-aware saliency prediction task, we adopt five commonly used consistency metrics: Area Under the Curve (AUC), Normalized Scanpath Saliency (NSS), Correlation Coefficient (CC), Similarity Metric (SIM), and Kullback–Leibler Divergence (KLD). For deep learning-based IQA and VQA models, we partition the Gen3DHF database into training and test sets at a same ratio of $4:1$ as the previous literature. The models are implemented with PyTorch and trained on a 40GB NVIDIA RTX A6000 GPU with batch size of 8. We use the Adam optimizer with the initial learning rate set as $41e{-}$ 4 and set the batch size as 4. The training process is stopped after 50 epochs. The resolution of input frames is rescale to 224 $\times$ 224. For deep-learning based saliency predict models, we retrain them on the proposed Gen3DHF database using their officially realeased code and utilizing the same training, testing sets as mention above. The training parameters remain the same, but is conducted for only 30 epochs with an early stopping strategy applied. All experiments for each LMME3DHF are averaged using 5-fold cross-validation.
5.2. Performance on Score Prediction and Visual Question Answering
@@
As shown in Table 1, traditional handcrafted IQA metrics such as BRISQUE (Mittal et al., 2012a) and NIQE (Mittal et al., 2012b) perform poorly in our scenario, indicating that their handcrafted features are primarily designed for natural image distortions and do not generalize well to AI-generated 3D human faces. On the other hand, although LLM-based metrics are widely recognized for their advanced visual understanding and visual question-answering capabilities, they fall short in accurately assessing perceptual video quality. In contrast, deep learning-based metrics, whether initially designed for IQA or VQA demonstrate significantly better performance compared to both handcrafted and LLM-based approaches. However, while these models achieve moderate to high performance in perceptual quality assessment, they generally lack visual question-answering capabilities, which are essential for improving the interpretability and diagnostic feedback in AI-generated content evaluation. Our proposed method LMME3DHF achieves the best performance from both quality and authenticity perspectives. This confirms the effectiveness of our model in evaluating human visual experience for AI-generated 3D HFs from multiple perspectives.
To assess the model performance on visual question answering, we further launch comparison of visual question-answering performance between our proposed LMME3DHF and various LMM-based metrics, as shown in Table 2. Models are asked to identify and classify the types of distortions present in the video content. They are required to select from predefined distortion categories as defined in Section 3. As the results show, LMME3DHF significantly outperforms all other baselines, achieving high accuracy in distortion type prediction, highlighting its strength in both perceptual understanding and detailed diagnostic capabilities.

### v3

@@ Abstract.
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3746027.3755700conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Ireland.booktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelandisbn: 979-8-4007-2035-2/2025/10ccs: Information systems Multimedia databasesccs: Information systems Multimedia streamingccs: Information systems Multimedia content creation
@@ 1. Introduction
In this paper, we introduce Gen3DHF, a comprehensive dataset and benchmark comprising 2,000 diverse 3D HF video samples generated by five distinct models. As illustrated in Figure 1, we collect 90,000 human annotations, resulting in 4,000 Mean Opinion Scores (MOS) and 2,000 distortion-aware saliency maps with corresponding distortion descriptions. Based on Gen3DHF, we propose LMME3DHF, a LMM-based metric specifically designed to not only evaluate 3D HF content across the two dimensions, i.e., quality and authenticity, but also to predict and output the salient distortion regions along with their corresponding textual descriptions. LMME3DHF leverages instruction tuning (Liu et al., 2023b) and LoRA adaptation(Hu et al., 2022) techniques to fine-tune the language model.
@@
  • Extensive experiments demonstrate LMME3DHF’s state-of-the-art performance in both accurately predicting quality scores for AI-generated 3D human faces and effectively identifying distortion-aware salient regions and distortion types.
@@ 2. RELATED WORK
Recent advancements in 3D-aware generative adversarial networks (GANs) significantly improve the quality of synthesized human faces, achieving breakthroughs in photorealism, multi-view consistency, and fine-grained controllability. While explicit and implicit neural rendering methods each have limitations in scalability or fidelity, (Chan et al., 2022) introduced a hybrid explicit–implicit tri-plane architecture that achieves high-resolution, multi-view-consistent synthesis from unstructured 2D images. Building on it, (An et al., 2023; Sun et al., 2023b; Wu et al., 2022b, 2023b; Deng et al., 2024b, a; Or-El et al., 2022; Schwarz et al., 2022; Zhang et al., 2022; Sun et al., 2022b) extend 3D HF generation to full $360^{\circ}$ head views, introduce animatable controls over facial expressions and head–shoulder movements respectively, detailed facial deformation with high fidelity, demonstrated high-fidelity 4D head synthesis and motion transfer using synthetic or pseudo multi-view data, achieving robust motion control without relying on monocular 3D Morphable Model (3DMM) reconstruction. Nevertheless, 3D human face generation still suffers from certain distortions, such as inconsistent views across frames, unrealistic geometry in unseen regions, or artifacts introduced by limited training data and view ambiguity.
@@ 2.2. Video Quality Assessment
With the rapid advancement of AI-generated image (AIGI), various image quality assessment (IQA) methods (Li et al., 2023b; Wang et al., 2023; Liang et al., 2024; Kirstain et al., 2023; Wu et al., 2023a) have emerged to evaluate the perceptual quality of these images. As the capabilities of AI-generated content continue to evolve (Dhariwal and Nichol, 2021; Rombach et al., 2022), AI-generated videos (AGVs) are now receiving increasing attention and being applied across a wide range of use cases. To assess their quality, several video quality assessment (VQA) studies have been conducted (Chen et al., 2024a; Chivileva et al., 2023; Liu et al., 2024, 2023a; Kou et al., 2024; Huang et al., 2024a). Early VQA methods often adopted evaluation metrics originally designed for IQA. While effective for measuring basic perceptual similarity, these metrics do not fully capture temporal coherence or motion consistency in videos. More recently, Contrastive Language–Image Pre-training (CLIP) (Radford et al., 2021) has been employed as a backbone for quality assessment tasks, leveraging its powerful multimodal representations to better align generated content with human-perceived semantics. Meanwhile, the rise of video large multimodal models introduces semantic-aware quality assessment, where models pre-trained on vision-language tasks jointly analyze visual fidelity and textual descriptions and particularly excel in visual question-answering. Existing deep learning-based VQA methods lack interactive QA capabilities, while video large multimodal models show limited precision in perceptual quality assessment.
@@ 2.3. Saliency Prediction
Recent advances in computer vision have driven significant progress in human visual attention analysis, with saliency prediction emerging as a crucial research area. This technology simulates human visual system’s mechanisms for allocating attention across different image regions. In recent years, numerous research efforts have focused on developing saliency prediction models. Traditional models (Bruce and Tsotsos, 2005; Goferman et al., 2011; Erdem and Erdem, 2013) primarily rely on low-level features such as intensity, color, and contrast to generate saliency maps. In contrast, deep learning-based models (Huang et al., 2015; Che et al., 2019; Lou et al., 2022) are capable of capturing more complex visual patterns and semantic information. However, despite their success in general applications, these models often fall short in the context of quality assessment, where identifying distortion-specific regions is crucial for accurately evaluating perceptual quality. This motivates us to develop a saliency prediction model specifically tailored for identifying distortion-aware salient regions in AI-generated human faces as distortion-aware saliency prediction not only guides model attention toward visually critical regions but also enhances interpretability by highlighting the exact areas that degrade user experience.
@@ 3.2. Subjective Experiment
We propose a dual-perspective evaluation framework to better align VQA with human preferences. The quality dimension assesses perceptual attributes such as texture, color fidelity, and structural integrity. In contrast, the authenticity dimension evaluates perceptual aspects like unidentified object appearance, head structure and the presence of unrealistic textures or shapes.
Subjective experiments are conducted in accordance with the ITU-R BT.500-13 (Series, 2012) guidelines. As shown in Figure 4, the interface presents a video along with a static image containing three snapshots from the video at $45^{\circ}$, $0^{\circ}$ and $45^{\circ}$ angles (with $0^{\circ}$ representing the frontal view). Participants can navigate the 3D HF projection videos using “Previous”, “Replay” and “Next” interactive buttons. Additionally, two sliders ranging from 0 to 5 enable participants to score the content on both quality and authenticity dimensions smoothly. A “MARK” button is provided to allow participants to highlight any visible distortions on the static images. If distortions are marked, participants are instructed to provide descriptions by typing or voice recording through “Voice” button. A total of 24 subjects participate in the subjective experiment. All participants receive a thorough briefing beforehand. Each session lasts approximately two hours. Finally, we obtain a total of 90,000 human annotations including 84,000 reliable score ratings (21 annotators $\times$ 2 dimensions $\times$ 2,000 videos), 6,000 distortion-aware saliency maps and distortion descriptions (3 annotators $\times$ 2,000 images) which were then aggregated into a single combined annotation per image.
  (a)              (b)
@@ 3.3. Subjective Data Processing
Additionally, to quantify performance and standardize raw distortion descriptions, we classify distortions into nine distinct categories based on their visual characteristics $d_{i}$, where $\{d_{i}|_{i=1}^{9}\}=$ {”Eye Distortions”, ”Mouth Distortions”, ”Hair Distortions”, ”Facial Feature Distortions”, ”Head Structure Distortions”, ”Overlap or Blending Issues”, ”Blurring / Exposure / Grain”, ”Accessories or Cloth Distortions”, “No Distortion”}. Due to the nature of our subjective experiment data where distortion regions are mark using red dots, a transformation is required to convert these discrete annotations into continuous distortion-aware saliency maps. As illustrated in Figure 5, and following standard practices in saliency prediction tasks, we first identify the center of each distorted region based on the human-marked red dots. These fixation maps are then smoothed using a Gaussian filter with a standard deviation of $\sigma=5$.
@@ 3.4. Subjective Data Analysis
The MOS and score distributions for the quality and authenticity dimensions are illustrated in Figure 2(a). Both distributions are slightly right-skewed, indicating that most samples received moderate to high ratings, peaking around the 55–60 range. This suggests that the majority of the generated 3D human face (HF) content was perceived as reasonably convincing in terms of both visual quality and authenticity. To further investigate the perceptual performance of different 3D HF generation models, we compare their MOS for quality and authenticity. As shown in Figure 2(b), EG3D (Chan et al., 2022) performs well in terms of quality but lags in authenticity, whereas PanoHead (An et al., 2023) exhibits the opposite trend. This contrast underscores the importance of evaluating quality and authenticity as separate dimensions. As illustrated in Figure 3(a), each distortion type occurs regularly, with facial feature distortions appearing most frequently. This indicates that AI-generated 3D HFs particular challenges in generating smooth, realistic, and detailed facial features. The distribution of distortion types across different generation models is illustrated in Figure 3(b). As shown in Figure 6, we observe that an increased number of distortion-aware salient regions strongly correlates with a significantly worse human visual experience. In other words, lower scores are associated with more extensive and concentrated saliency in distorted areas.
Figure 7. Overview of the LMME3DHF architecture. The model includes three functions: (a) text-defined quality level prediction, (b) precises score prediction, (c) distortion-aware saliency map prediction. The training process consists of three stages: (a) instruction tuning of the model via text-defined levels and distortion description, (b) fine-tuning the LMM via numerical scores, (c) train saliency decoder from scratch.
@@ 4.1. Model Structure
Visual Encoding.  As shown in Figure 7, we leverage the pretrained Vision Transformer (ViT) module embedded in Qwen2.5-VL-7B to extract visual features. This model demonstrates strong capabilities in handling videos with varying resolutions and frame rates, allowing us to bypass preprocessing steps such as compression or resizing, which often lead to information loss. The native projector, implemented using two-layer Multi-Layer Perceptron (MLP), is used to project the extracted visual features into a latent space aligned with the text embeddings used in the language model.
@@ 4.2. Training and Fine-tuning Strategy
Instruction Tuning.  Due to the inherent strengths of LLMs in processing textual rather than numerical data, we begin by converting quality scores into categorical, text-based quality levels. We uniformly divides the score range from the minimum value (m) to the maximum value (M) into five equal intervals and assigns each interval a corresponding textual label based on Equation (4).
@@
where $\{l_{i}|_{i=1}^{5}\}=\{bad,poor,fair,good,excellent\}$. The distortion descriptions were categorized into nine classes, as previous literature described. During instruction tuning, the model is trained using language loss, guiding it to generate quality-level predictions and distortion category classifications.
Quality Regression Fine-Tuning.  To enable continuous quality score prediction, we implement a quality regression module. This module takes the last hidden states from the LLM as input and outputs precise predicted scores. As full fine-tuning of LLMs can be computationally expensive, we leverage the LoRA (Hu et al., 2022) technique to efficiently adapt the model to the quality regression task. During this fine-tuning phase, we leverage L1 loss to align predicted scores with ground-truth MOS.
@@ 5.1. Experiment Setup
To evaluate the correlation between predicted scores and the ground-truth MOS, we employ three evaluation metrics: Spearman’s Rank Correlation Coefficient (SRCC) (Spearman, 1904), Pearson’s Linear Correlation Coefficient (PLCC) (Pearson, 1896), and Kendall’s Rank Correlation Coefficient (KRCC) (Kendall, 1938). For assessing visual question answering performance, we use average accuracy as the evaluation metric. For the distortion-aware saliency prediction task, we adopt five commonly used consistency metrics: Area Under the Curve (AUC) (Hanley and McNeil, 1982), Normalized Scanpath Saliency (NSS) (Peters et al., 2005), Correlation Coefficient (CC) (Pearson, 1896), Similarity Metric (SIM) (Swain and Ballard, 1991), and Kullback–Leibler Divergence (KLD) (Kullback and Leibler, 1951). For deep learning-based IQA and VQA models, we partition the Gen3DHF database into training and test sets at a same ratio of $4:1$ as the previous literature. The models are implemented with PyTorch and trained on a 40GB NVIDIA RTX A6000 GPU with batch size of 8. We use the Adam optimizer with the initial learning rate set as 1e41e{-}4 and set the batch size as 4. The training process is stopped after 50 epochs. For deep-learning based saliency predict models, we retrain them on the proposed Gen3DHF database using their officially realeased code and utilizing the same training, testing sets as mention above. The training parameters remain the same, but is conducted for only 30 epochs with an early stopping strategy applied. All experiments for each LMME3DHF are averaged using 5-fold cross-validation.
@@
5.2. Performance on Score Prediction and Visual Question Answering
As shown in Table 1, traditional handcrafted IQA metrics such as BRISQUE (Mittal et al., 2012a) and NIQE (Mittal et al., 2012b) perform poorly in our scenario, indicating that their handcrafted features are primarily designed for natural image distortions and do not generalize well to AI-generated 3D human faces. On the other hand, although LLM-based metrics are widely recognized for their advanced visual understanding and visual question-answering capabilities, they fall short in accurately assessing perceptual video quality. In contrast, deep learning-based metrics, whether initially designed for IQA or VQA demonstrate significantly better performance compared to both handcrafted and LMM-based approaches. However, while these models achieve moderate to high performance in perceptual quality assessment, they generally lack visual question-answering capabilities, which are essential for improving the interpretability and diagnostic feedback in AI-generated content evaluation. Our proposed method LMME3DHF achieves the best performance from both quality and authenticity perspectives. This confirms the effectiveness of our model in evaluating human visual experience for AI-generated 3D HFs from multiple perspectives.
To assess the model performance on visual question answering, we further launch comparison of visual question-answering performance between our proposed LMME3DHF and various LMM-based metrics, as shown in Table 2. Models are asked to identify and classify the types of distortions present in the video content. As the results show, LMME3DHF significantly outperforms all other baselines, highlighting its strength in both perceptual understanding and detailed diagnostic capabilities.